# AGN - host galaxy photometric decomposition using a fast, accurate and precise deep learning approach

### v1

@@ Abstract
Aims. Identification of active galactic nuclei (AGNs) is extremely important for understanding galaxy evolution and its connection with the formation and evolution of supermassive black holes (SMBH). With the advent of deep and high angular resolution imaging surveys such as those conducted with the James Webb Space Telescope (JWST), it is now possible to identify galaxies with a central point source out to the very early Universe. In this proof of concept study, we aim to develop a fast, accurate and precise method to identify galaxies which host AGNs and recover the intrinsic AGN contribution to the observed total light ($f_{AGN}$).
Methods. We trained a deep learning (DL) based method Zoobot to estimate the fractional contribution of a central point source to the total light. Our training sample comprises realistic mock JWST images of simulated galaxies from the IllustrisTNG cosmological hydrodynamical simulations. We injected different amounts of the real JWST point spread function (PSF) models to represent galaxies with different levels of AGN contribution. Galaxies in our training sample span a wide range of morphologies, including merging galaxies. We analyse in detail the performance of our method as a function of various galaxy properties and compare it with results obtained from the traditional light profile fitting tool GALFIT.
Results. We find excellent performance of our DL method in recovering the injected AGN fraction $f_{AGN}$, both in terms of precision and accuracy. The mean difference between the predicted and true injected $f_{AGN}$ is $-0.0006$ and the overall root mean square error (RMSE) is $0.027$. The relative absolute error (RAE) is $0.086$ and the outlier (defined as predictions with RAE $>20\%$) fraction is $7.8\%$. In comparison, using GALFIT on the same dataset, we achieve a mean difference of -$0.024$, RMSE of $0.09$, RAE of $0.19$ and outlier fraction of $19\%$. We also investigate quantitatively how these key performance metrics obtained from Zoobot and GALFIT vary as a function of the injected $f_{AGN}$, redshift, signal-to-noise ratio, and magnitude. In addition to superior performance, our DL method has several other advantages over traditional methods, e.g. it has a much higher success rate (even for highly disturbed or irregular galaxies) and is extremely fast and easy to use. In summary, our DL-based method to identify AGNs and estimate AGN contribution to the total light has a huge potential in future applications to large galaxy imaging surveys.
@@
Galaxies: interactions – Galaxies: active – Galaxies: evolution – Techniques: image processing
@@ 1 Introduction
It is generally accepted that there exists co-evolution between supermassive black holes (SMBH) and their host galaxies (Kormendy & Ho, 2013). One manifestation of this co-evolution is the tight correlation between the mass of an SMBH and its host galaxy properties, such as the stellar velocity dispersion, bulge luminosity and bulge mass (Gültekin et al., 2009; Beifiori et al., 2012; Graham & Scott, 2013; McConnell & Ma, 2013; Läsker et al., 2014). These correlations may hint at a fundamental connection between the central SMBH and the formation and evolution of its host galaxy. Theoretical models show that SMBH feedback, via radiative heating, outflows, or jets could potentially explain these correlations, by regulating or halting the growth of itself and the host galaxy (e.g., Somerville et al., 2008; Booth & Schaye, 2009; Weinberger et al., 2018; Davé et al., 2019). However, these correlations could also arise from merging events, in which mergers could explain, for example, the growth of SMBH mass and the stellar mass of the host galaxy (e.g., Croton, 2006; Peng, 2007; Hirschmann et al., 2010; Jahnke & Macciò, 2011). To better understand the co-evolution (or not) of SMBHs and host galaxies, it is important to study their link at different cosmic times. At high redshifts, studies have to rely on accreting SMBHs, which are active galactic nuclei (AGNs), to be able to obtain SMBH mass estimates, and in particular Type I AGNs. It is also necessary to obtain quantitative measures of the physical properties of the host galaxies. However, the presence of a bright central AGN can make this task very difficult, particularly for the bulge component, as galaxies with significant contribution from the AGN to the total flux will appear more bulge-dominated (Pierce et al., 2010).
@@
It is crucial to correctly separate the central AGN light from the host galaxy, across a wide range of galaxy types and redshifts. Good quality optical imaging, with high spatial resolution and signal-to-noise, is needed to decompose the observed total light into contributions from the host galaxy and the central AGN. Traditionally, this is done in photometric data by fitting two-dimensional (2D) profiles to the galaxy’s light, using one (or more if needed) analytic profile to describe the galaxy (typically a Sérsic profile) and a point spread function (PSF) profile to describe the central point source. Many studies use GALFIT (Peng et al., 2002), one of the most widely used 2D surface brightness modelling software to perform image-based decomposition of AGN and host galaxy light (e.g. Kim et al., 2008; Bentz et al., 2009; Gabor et al., 2009; Böhm et al., 2013; Schramm & Silverman, 2013; Du et al., 2014; Urbano-Mayorgas et al., 2019; Son et al., 2022; Aird et al., 2022; Ji et al., 2022; Dewsnap et al., 2023; Zhuang & Ho, 2023; Sturm & Reines, 2024; Zhuang et al., 2024). However, there are some known technical issues with GALFIT. For example, in certain circumstances, particularly for extremely compact sources, the minimisation algorithm used can be trapped in a local minimum. Other software have been used in an attempt to improve on GALFIT performance, such as PSFMC (Mechtley, 2014), which is a Markov Chain Monte Carlo (MCMC) simultaneous fitting software to perform multi-component profile fitting (Mechtley et al., 2016; Marshall et al., 2021) and LENSTRONOMY (Birrer et al., 2015; Birrer & Amara, 2018), which uses particle swarm optimisation (Kennedy & Eberhart, 1995) for $\chi^{2}$ minimisation, to reduce the likelihood of getting trapped in a local minimum when searching the parameter space, and MCMC for Bayesian parameter inference (Foreman-Mackey et al., 2013). The latter method is used by (Li et al., 2021) to decompose a sample of X-ray selected AGNs into quasars and host galaxy components, to investigate the properties of the host galaxies. However, these approaches all assume that the galaxy’s surface brightness profile can be well fitted by a single Sérsic profile (or combination of Sérsic profiles), which may not be the case, particularly for irregular galaxies, highly disturbed merging galaxies, or galaxies with complicated substructures.
@@
Consequently, if Sérsic profiles cannot adequately describe the host galaxy light, it can introduce a systematic bias in the derived luminosity of the PSF component and of the host galaxy. In addition, if there is a significant contribution to the total flux from the central AGN, it can lead to a bias in the morphology of the host galaxy, as it may appear more bulge-dominated (Pierce et al., 2010). Indeed, it is not always a good idea to represent a galaxy with a single Sérsic profile. Bentz et al. (2009) study AGN host galaxies at redshift $z\approx 0.7$, by fitting their surface brightness distributions with a combination of Sérsic profiles (to account for the host galaxy’s light) and PSF profile (to describe the AGN component). They find that most galaxies in their sample are well described with two Sérsic profiles, in addition to the PSF profile, and a few of them required three or more Sérsic profiles to describe additional components such as bars. Another possible difficulty in decomposing the AGN from the host galaxy is the (sometimes significant) variation of the PSF in a given survey due to spatial/temporal changes or differences in galaxy spectral energy distributions (SEDs), as an incorrect PSF can bias the estimated contribution of the AGN to the total flux of the galaxy. Kim et al. (2008) perform 2D decomposition of Sérsic + PSF profiles to galaxies hosting AGNs in Hubble Space Telescope (HST) images and investigates the effect of realistic PSF mismatch, finding a systematic overestimation of the flux of the host galaxies, particularly for those containing bright AGNs.
@@
Our DL model is trained on mock images of simulated galaxies with different levels of AGN contributions to the total flux. The construction of our training sample can easily incorporate the full information on the expected variations of the PSF. Therefore, our DL model can learn to infer the intrinsic AGN fraction automatically folding in the impact of different PSFs. In comparison, while it is possible to examine the goodness-of-fit for different PSF models for methods based on light profile fitting such as GALFIT, in practice it will be extremely time-consuming, particularly for large samples of galaxies. Another advantage of our DL-based method is that it does not rely on an assumed (and often simplified) galaxy surface brightness profile which in many cases is not able to fully describe a galaxy’s light profile and thus can introduce biases in the estimation of the AGN contribution. Finally, our method, like any other machine learning-based method, has the advantage of being very fast to implement in new data once the DL model has been trained, making it much more computationally efficient than any traditional method based on light profile fitting.
This paper is organised as follows. In Section 2, we describe the observed James Webb Space Telescope (JWST; Gardner et al., 2006) imaging data used in this work and the generation of the corresponding mock JWST images of simulated galaxies selected from the IllustrisTNG. Of particular importance are the real JWST PSF models and their variations within the survey data. In Section 3, first we explain how we create the final images mimicking galaxies containing AGNs by injecting different levels of PSF contribution (taken as AGN contribution) to mock JWST images. Then we introduce our DL-based method (Zoobot; Walmsley et al., 2023) to recover the intrinsic AGN fraction in the observed total light. To compare with traditional light profile fitting-based methods, we also briefly describe GALFIT and how we use it in this work. In Section 4, we present in detail the results from both methods and compare their performances as a function of various galaxy properties. Finally, in Section 5 we summarise the paper and highlight the main conclusions of our work. Throughout the paper we assume a flat $\Lambda$CDM universe with $\Omega_{M}=0.2865$, $\Omega_{\Lambda}=0.7135$, and $H_{0}=69.32$ km s-1 Mpc-1 (Hinshaw et al., 2013). Unless otherwise stated, magnitudes are presented in the AB system.
@@ 2 Data
In this section, we first introduce the real JWST observations and the PSF models obtained from the COSMOS-Web survey. Then, we describe the synthetic JWST images generated using simulated galaxies selected from the IllustrisTNG cosmological hydrodynamical simulations. The observed and simulated datasets are combined later to create mock galaxy JWST images with different levels of AGN contribution.
@@ 2.1 JWST/COSMOS-WEB
COSMOS-Web (Casey et al., 2023, PIs: Kartaltepe & Casey, ID=1727) is a 255-hour JWST treasury imaging survey observing the central area of the Cosmic Evolution Survey (COSMOS; Scoville et al., 2007) field which is one of the most popular deep multi-wavelength survey fields. It covers a contiguous 0.54 deg2 region in four Near Infrared Camera (NIRCam) filters (F115W, F150W, F277W, and F [^4] W), reaching a $5\sigma$ point source depth of $27.5-28.2$ magnitudes. In parallel, a 0.19 deg2 area of Mid-Infrared Instrument (MIRI) imaging with the F770W filter is covered. For this work, we make use of the 0.28 deg2 JWST/NIRCam F150W images reduced by Zhuang et al. (2024). They used version 1.10.2 of the JWST [^1] https://jwst-pipeline.readthedocs.io/en/latest/ pipeline with the Calibration Reference Data System (CRDS) version of 11.17.0. to reduce the uncalibrated NIRCam raw data retrieved from MAST [^2] https://archive.stsci.edu/doi/resolve/resolve.html?doi=10.17909/6btv-br09. The steps followed in the data reduction process are briefly summarised in the following:
@@
Figure 1: Overview of the PSFs. We stacked all available PSFs from Zhuang et al. (2024) and show the mean PSF (top panel), standard deviation (central panel) and the coefficient of variation (bottom panel), calculated pixel by pixel. The pixel resolution is $0.015\ \arcsec/\text{pixel}$. The axes show the number of pixels, corresponding to 3″. The colorbar shows the intensity of each pixel.

### v2

@@ Abstract
Aims. Identification of active galactic nuclei (AGN) is extremely important for understanding galaxy evolution and its connection with the assembly of supermassive black holes (SMBH). With the advent of deep and high angular resolution imaging surveys such as those conducted with the James Webb Space Telescope (JWST), it is now possible to identify galaxies with a central point source out to the very early Universe. In this proof of concept study, we aim to develop a fast, accurate and precise method to identify galaxies which host AGN and recover the intrinsic AGN contribution fraction ($f_{AGN}$).
Methods. We trained a deep learning (DL) based method Zoobot to estimate the fractional contribution of a central point source to the total light. Our training sample comprises realistic mock JWST images of simulated galaxies from the IllustrisTNG cosmological hydrodynamical simulations. We injected different amounts of the observed JWST point spread functions to represent galaxies with varying levels of AGN contribution. Galaxies in our training sample span a wide range of morphologies, including mergers. We analyse in detail the performance of our method as a function of various galaxy properties and compare with results obtained from the traditional light profile fitting tool GALFIT. After training, we applied our method to real JWST observations.
Results. We find excellent performance of our DL method in recovering the injected $f_{AGN}$, in terms of precision and accuracy. The mean difference between the predicted and true $f_{AGN}$ is $-0.002$ and the overall root mean square error (RMSE) is $0.013$. The overall relative absolute error (RAE) is $0.076$ and the outlier (defined as predictions with RAE $>20\%$) fraction is $6.5\%$. In comparison, using GALFIT, we achieve a mean difference of -$0.02$, RMSE of $0.12$, RAE of $0.19$ and outlier fraction of $19\%$. We also investigate how these key performance metrics obtained from Zoobot and GALFIT vary as a function of the injected $f_{AGN}$, redshift, signal-to-noise ratio, and galaxy size. In addition to the superior performance, our DL method has several other advantages over traditional methods. For example, it has a much higher success rate (even for highly disturbed or irregular galaxies) and is extremely fast. We applied our trained DL model to real JWST observations and found that 20% of the X-ray-selected AGN and 8% of the MIR-selected AGN are also identified as AGN using a cut at $f_{\rm AGN}>0.2$. When using $f_{\rm AGN}>0.1$, these overlaps increase to 33% for the X-ray AGN and 15% for the MIR AGN. In summary, our DL-based method to identify AGN and estimate AGN contribution fraction has a huge potential in future applications to large galaxy imaging surveys.
@@
Galaxies: active – Galaxies: evolution – Techniques: image processing
@@ 1 Introduction
It is generally accepted that there exists co-evolution between supermassive black holes (SMBH) and their host galaxies (Kormendy & Ho, 2013). One manifestation of this co-evolution is the tight correlation between the mass of an SMBH and its host galaxy properties, such as the stellar velocity dispersion, bulge luminosity and bulge mass (Gültekin et al., 2009; Beifiori et al., 2012; Graham & Scott, 2013; McConnell & Ma, 2013; Läsker et al., 2014). These correlations may hint at a fundamental connection between the central SMBH and the formation and evolution of its host galaxy. Theoretical models show that SMBH feedback, via radiative heating, outflows, or jets could potentially explain these correlations, by regulating or halting the growth of itself and the host galaxy (e.g., Somerville et al., 2008; Booth & Schaye, 2009; Weinberger et al., 2018; Davé et al., 2019). However, these correlations could also arise from merging events, in which mergers could explain, for example, the growth of SMBH mass and the stellar mass of the host galaxy (e.g., Croton, 2006; Peng, 2007; Hirschmann et al., 2010; Jahnke & Macciò, 2011). To better understand the co-evolution (or not) of SMBHs and host galaxies, it is important to study their link at different cosmic times. At high redshifts, studies have to rely on accreting SMBHs, which are active galactic nuclei (AGN), to be able to obtain SMBH mass estimates, and in particular Type I AGN. It is also necessary to obtain quantitative measures of the physical properties of the host galaxies. However, the presence of a bright central AGN can make this task very difficult, particularly for the bulge component, as galaxies with significant contribution from the AGN to the total flux will appear more bulge-dominated (Pierce et al., 2010).
@@
It is crucial to correctly separate the central AGN light from the host galaxy, across a wide range of galaxy types and redshifts. Good quality optical imaging, with high spatial resolution and signal-to-noise ratio (S/N), is needed to decompose the observed total light into contributions from the host galaxy and the central AGN. Traditionally, this is done in photometric data by fitting two-dimensional (2D) profiles to the galaxy’s light, using one (or more if needed) analytic profile to describe the galaxy (typically a Sérsic profile) and a point spread function (PSF) profile to describe the central point source. Many studies use GALFIT (Peng et al., 2002), one of the most widely used 2D surface brightness modelling software to perform image-based decomposition of AGN and host galaxy light (e.g. Kim et al., 2008; Bentz et al., 2009; Gabor et al., 2009; Böhm et al., 2013; Schramm & Silverman, 2013; Du et al., 2014; Urbano-Mayorgas et al., 2019; Son et al., 2022; Aird et al., 2022; Ji et al., 2022; Dewsnap et al., 2023; Zhuang & Ho, 2023; Sturm & Reines, 2024; Zhuang et al., 2024). However, there are some known technical issues with GALFIT. For example, in certain circumstances, particularly for extremely compact sources, the minimisation algorithm used can be trapped in a local minimum. Other software have been used in an attempt to improve on GALFIT performance, such as PSFMC (Mechtley, 2014), which is a Markov Chain Monte Carlo (MCMC) simultaneous fitting software to perform multi-component profile fitting (Mechtley et al., 2016; Marshall et al., 2021) and LENSTRONOMY (Birrer et al., 2015; Birrer & Amara, 2018), which uses particle swarm optimisation (Kennedy & Eberhart, 1995) for $\chi^{2}$ minimisation, to reduce the likelihood of getting trapped in a local minimum when searching the parameter space, and MCMC for Bayesian parameter inference (Foreman-Mackey et al., 2013). The latter method is used by (Li et al., 2021) to decompose a sample of X-ray-selected AGN into quasars and host galaxy components, to investigate the properties of the host galaxies. However, these approaches all assume that the galaxy’s surface brightness profile can be well fitted by a single Sérsic profile (or combination of Sérsic profiles), which may not be the case, particularly for irregular galaxies, highly disturbed merging galaxies, or galaxies with complicated substructures.
@@
Consequently, if Sérsic profiles cannot adequately describe the host galaxy light, it can introduce a systematic bias in the derived luminosity of the PSF component and of the host galaxy. In addition, if there is a significant contribution to the total flux from the central AGN, it can lead to a bias in the morphology of the host galaxy, as it may appear more bulge-dominated (Pierce et al., 2010). Indeed, it is not always a good idea to represent a galaxy with a single Sérsic profile. Bentz et al. (2009) study AGN host galaxies at redshift $z\approx 0.7$, by fitting their surface brightness distributions with a combination of Sérsic profiles (to account for the host galaxy’s light) and PSF profile (to describe the AGN component). They find that most galaxies in their sample are well described with two Sérsic profiles, in addition to the PSF profile, and a few of them required three or more Sérsic profiles to describe additional components such as bars. Another possible difficulty in decomposing the AGN from the host galaxy is the (sometimes significant) variation of the PSF in a given survey due to spatial/temporal changes or differences in galaxy spectral energy distributions (SEDs), as an incorrect PSF can bias the estimated contribution of the AGN to the total flux of the galaxy. Kim et al. (2008) perform 2D decomposition of Sérsic + PSF profiles to galaxies hosting AGN in Hubble Space Telescope (HST) images and investigates the effect of realistic PSF mismatch, finding a systematic overestimation of the flux of the host galaxies, particularly for those containing bright AGN.
@@
Our DL model is trained on mock images of simulated galaxies with different levels of AGN contributions to the total flux. The construction of our training sample can easily incorporate the full information on the expected variations of the PSF. Therefore, our DL model can learn to infer the intrinsic AGN fraction while automatically folding in the impact of different PSFs. In comparison, while it is possible to examine the goodness-of-fit for different PSF models for methods based on light profile fitting such as GALFIT, in practice it will be extremely time-consuming, particularly for large samples of galaxies. Another advantage of our DL-based method is that it does not rely on an assumed (and often simplified) galaxy surface brightness profile which in many cases is not able to fully describe a galaxy’s light profile and thus can introduce biases in the estimation of the AGN contribution. Finally, our method, like any other machine learning-based method, has the advantage of being very fast to implement in new data once the DL model has been trained, making it much more computationally efficient than any traditional method based on light profile fitting.
This paper is organised as follows. In Section 2, we describe the observed James Webb Space Telescope (JWST; Gardner et al., 2006) imaging data used in this work and the generation of the corresponding mock JWST images of simulated galaxies selected from the IllustrisTNG. Of particular importance are the real JWST PSF models and their variations within the survey data. In addition, we introduce two AGN samples, selected in the X-ray and mid-infrared (MIR), which we use to compare with AGN identified using our DL-based method. In Section 3, first we explain how we create the final images mimicking galaxies containing AGN by injecting different levels of PSF contribution (taken as AGN contribution fractions) in the mock JWST images. Then we introduce our DL-based method (Zoobot; Walmsley et al., 2023) to recover the intrinsic AGN contribution fraction in the observed total light. To compare with traditional light profile fitting-based methods, we also briefly describe GALFIT and how we use it in this work. In Section 4, we explore in detail the results from both methods and compare their performances as a function of various galaxy properties. We also present the first application of our DL-based method to real JWST data and compare with AGN selected in the X-ray and MIR. Finally, in Section 5 we summarise the paper and highlight the main conclusions of our work.
Throughout the paper we assume a flat $\Lambda$CDM universe with $\Omega_{M}=0.2865$, $\Omega_{\Lambda}=0.7135$, and $H_{0}=69.32$ km s-1 Mpc-1 (Hinshaw et al., 2013). Unless otherwise stated, magnitudes are presented in the AB system.
@@ 2 Data
In this section, we first introduce the real JWST observations and the PSF models obtained from the COSMOS-Web survey. Then, we describe the synthetic JWST images generated using simulated galaxies selected from the IllustrisTNG cosmological hydrodynamical simulations. The observed and simulated datasets are combined later to create mock JWST galaxy images with different levels of AGN contribution.
@@ 2.1 JWST/COSMOS-WEB
Figure 1: Overview of the JWST/NIRCAM F150W PSFs in COSMOS-Web. We stacked all available PSFs from Zhuang et al. (2024) and show the mean PSF (left), standard deviation (centre) and the coefficient of variation (right), calculated pixel by pixel. The pixel resolution is $0.015\ \arcsec/\text{pixel}$. The axes show the number of pixels, corresponding to 3″across. The colorbar shows the value of each pixel.
COSMOS-Web (Casey et al., 2023, PIs: Kartaltepe & Casey, ID=1727) is a 255-hour JWST treasury imaging survey observing the central area of the Cosmic Evolution Survey (COSMOS; Scoville et al., 2007) field which is one of the most popular deep multi-wavelength survey fields. It covers a contiguous 0.54 deg2 region in four Near Infrared Camera (NIRCam) filters (F115W, F150W, F277W, and F [^4] W), reaching a $5\sigma$ point source depth of $27.5-28.2$ magnitudes. In parallel, a 0.19 deg2 area of Mid-Infrared Instrument (MIRI) imaging with the F770W filter is covered. For this work, we make use of the 0.28 deg2 JWST/NIRCam F150W images reduced by Zhuang et al. (2024). They used version 1.10.2 of the JWST [^1] https://jwst-pipeline.readthedocs.io/en/latest/ pipeline with the Calibration Reference Data System (CRDS) version of 11.17.0. to reduce the uncalibrated NIRCam raw data retrieved from MAST [^2] https://archive.stsci.edu/doi/resolve/resolve.html?doi=10.17909/6btv-br09. The steps followed in the data reduction process are briefly summarised below:
@@
We use the COSMOS2020 (Weaver et al., 2022) photometric catalogue to construct our sample of real JWST galaxies, within the COSMOS-Web area. For that, we use the Farmer’s version for the photometric catalogue. We use spectroscopic redshifts when available. Otherwise, we adopt the photometric redshifts (photo-$z$) listed in the COSMOS2020 catalogue, computed using the LePhare code (Arnouts et al., 2002; Ilbert et al., 2006). The photo-$z$ precision, given by the normalised median absolute deviation, is around $0.01\times(1+z)$ at $i<24.0$ mag, and $0.03\times(1+z)$ at $27.024.0<i<$ 27.0 mag. Stellar masses ($M_{*}$) are derived by La Marca et al. 2025 (in prep.), using the spectral energy distribution fitting tool CIGALE (Burgarella et al., 2005; Noll et al., 2009; Boquien et al., 2019), including AGN models. These stellar masses are consistent with the stellar masses presented in the COSMOS2020 photometric catalogue obtained by LePhare. We calculate the median bias, $b$, and the median absolute deviation (MAD) between the two stellar mass estimates as follows:
$b=\text{median}(\Delta M_{*}),$ (1)
$MAD=1.48\times\text{median}(|\Delta M_{*}-\text{median}(\Delta M_{*})|)\,,$ (2)
where $\Delta M_{*}=log_{10}M_{*,LePhare}-log_{10}M_{*,CIGALE}$. We find $b=-0.09$ and $MAD=0.013$. We construct our galaxy sample in the redshift range $30.5<z<$ 3, avoiding stars or areas affected by bright stars. Finally, we select a stellar mass complete sample, by including galaxies with $M_{*}$ greater than either $10^{9}\,M_{\odot}$ or the KS-based completeness limit from Weaver et al. (2022) for the COSMOS2020 sample, whichever is higher (the minimum stellar mass limit of $10^{9}\,M_{\odot}$ is imposed to match the limit in the simulated galaxies). In total, we include 25 596 galaxies in our stellar mass complete sample over the 0.28 deg2 analysed in this study.
2.2 AGN selections
We use the following two common AGN selection techniques to identify AGN in our JWST sample and compare with our DL-based methodology in Sect. 4.3,
  1. 1.
    X-ray AGN selection: For this selection, we use the X-ray photometry data from the Chandra COSMOS Legacy survey (Civano et al., 2016; Marchesi et al., 2016). From the catalogue of optical and infrared counterparts provided by Marchesi et al. (2016), we select sources with a final counterpart identification flag of 1 (secure) or 10 (ambiguous). We then cross-matched this catalogue using the optical coordinates with our selected sources in COSMOS2020. We selected secure X-ray AGN requiring det_ml (the maximum likelihood detection) ¿10.8 in the hard or soft band.
  2. 2.
    MIR AGN selection: For this selection, we use the Spitzer/MIPS $24\mu m$ data, provided by the COSMOS-Spitzer programme (Sanders et al., 2007), and the MIR data from the four channels of the Spitzer/IRAC Cosmic Dawn Survey (Euclid Collaboration et al. 2022). We follow Chang et al. (2017) to identify AGN by their MIR emission. We limited the selection to those galaxies with Spitzer/MIPS 24 $\mu$m flux $F_{24\mu m}>20\,\mu$Jy, which is the $1\sigma$ total noise (instrument and confusion noise). Additionally, we required the signal-to-noise ratio ($S/N>5$) in each IRAC channel. Then, we applied the colour--colour criteria in Chang et al. (2017) to select AGN:
    $y$ $<2.22\times x+1.01,$ (3)
    $y$ $<8.67\times x-0.28,$ (4)
    $y$ $>-0.33\times x+0.17,$ (5)
    $y$ $>0.31\times x-0.06,$ (6)
    where $x=m_{3.6\mu m}-m_{5.8\mu m}$ and $y=m_{4.5\mu m}-m_{8\mu m}$. Magnitudes are expressed as AB magnitude.